OpenStudy (anonymous):

what steady annual rate of inflation would cause the value of an item to rise from $2500 to $18400 in 25 years?

OpenStudy (anonymous):

yeah i know i need to solve that, but i don't know how too :/ @ganeshie8

ganeshie8 (ganeshie8):

Alright, we use compound interest formula \(A = P(1+r)^t\)

ganeshie8 (ganeshie8):

\(P\) = starting value = 2500 \(r\) = rate of growth \(t\) = time = 25

ganeshie8 (ganeshie8):

plug them and solve \(r\) : \(\large 18400 = 2500(1+r)^{25}\)

ganeshie8 (ganeshie8):

start by dividing 2500 both sides

OpenStudy (anonymous):

how do you do that @ganeshie8

ganeshie8 (ganeshie8):

\(\large 18400 = 2500(1+r)^{25}\) divide 2500 both sides, you get : \(\large \dfrac{18400}{2500} = (1+r)^{25}\) which is same as : \(\large \dfrac{184}{25} = (1+r)^{25}\)

ganeshie8 (ganeshie8):

fine, so far ?

OpenStudy (anonymous):

yeah, i did that @ganeshie8

ganeshie8 (ganeshie8):

good, next take `25th root` both sides

ganeshie8 (ganeshie8):

\(\large 18400 = 2500(1+r)^{25}\) divide 2500 both sides, you get : \(\large \dfrac{18400}{2500} = (1+r)^{25}\) which is same as : \(\large \dfrac{184}{25} = (1+r)^{25}\) take 25th root both sides : \(\large \left(\dfrac{184}{25}\right)^{\frac{1}{25}} = (1+r)\)

ganeshie8 (ganeshie8):

lastly subtract 1 both sides

ganeshie8 (ganeshie8):

\(\large 18400 = 2500(1+r)^{25}\) divide 2500 both sides, you get : \(\large \dfrac{18400}{2500} = (1+r)^{25}\) which is same as : \(\large \dfrac{184}{25} = (1+r)^{25}\) take 25th root both sides : \(\large \left(\dfrac{184}{25}\right)^{\frac{1}{25}} = (1+r)\) subtract 1 both sides : \(\large \left(\dfrac{184}{25}\right)^{\frac{1}{25}}-1 = r\)

ganeshie8 (ganeshie8):

plug the left chunk into ur calculator

ganeshie8 (ganeshie8):

you get 0.083, which is same as 8.3%

OpenStudy (anonymous):

okay, thank you so much!! @ganeshie8

ganeshie8 (ganeshie8):

you're welcome !